# Establishing a Foundation for Tetun Text Ad-Hoc Retrieval: Stemming, Indexing, Retrieval, and Ranking

### v2

@@ 2.4. Summary
In ad-hoc text retrieval, preprocessing steps, such as stopword removal and stemming, are often employed to enhance retrieval efficiency and effectiveness. However, studies show that the impact of these techniques on retrieval effectiveness varies across languages, proving beneficial in some instances but less so in others. Furthermore, Ghosh and Bhattacharya (Ghosh and Bhattacharya, 2017) highlight that the influence of stopwords can differ even within the same language across different collections, such as Bangla and Hindi in the FIRE datasets of 2010 and 2011.
@@
This study addresses a critical gap in Tetun text retrieval by introducing three essential resources: a stopword list (Labadain-Stopwords), a language-specific stemmer (Labadain-Stemmer), and a Tetun test collection (Labadain-Avaliadór). Through a series of experiments, we evaluate various retrieval strategies to establish baselines and identify the most effective approach for ad-hoc text retrieval in Tetun. The subsequent sections provide a detailed overview of the development of each resource and their application in the experiments.
@@ 3. Tetun
This section provides an overview of Tetun, including its orthography, morphology, and loanwords.
@@ 3.1. Overview
Tetun, alternatively written as Tetum (English) or Tétum (Portuguese), is an Austronesian language spoken in Timor-Leste, an island country in Southeast Asia. There are two primary varieties of Tetun: Tetun Dili, also known as Tetun Prasa (commonly referred to as Tetun), and Tetun Terik (van Klinken et al., 2002). Tetun has two standardization forms: the standard produced by the Instituto Nacional de Linguística (Tetun INL) and the standard of the Dili Institute of Technology (Tetun DIT), while Tetun Terik is one of the dialects spoken in Timor-Leste.
Tetun is one of Timor-Leste’s official languages alongside Portuguese (Vasconcelos et al., 2011). The government established Tetun INL as the standard orthography, and this standardized form is used in the education system, official publications, and media (DL 01/2004, 2004). Tetun DIT was developed by linguists at the Dili Institute of Technology with some standardized differences from Tetun INL in terms of writing conventions (van Klinken et al., 2002). For example, the words [ foo (give), neebe (which/that) ] in Tetun DIT correspond to [ ne’ebé ] in Tetun INL.
According to the 2015 census report, Timor-Leste’s population was 1.18 million, with 78.78% of the population being Tetun speakers (de Jesus and Nunes, 2024a). Among them, 30.50% considered Tetun their home language, while 48.28% spoke it as a second or third language. Census 2023 reported a population growth of 13.40%, increasing from 1.18 million to 1.34 million (INETL, 2022), but it did not provide specific indicators for Tetun speakers.
@@ 3.2. Orthography
Tetun INL is based on the Latin alphabet, consisting of 5 vowels: a, e, i, o, u, and 21 consonants: b, d, f, g, h, j, k, l, ll, m, n, ñ, p, r, rr, s, t, ’, v, x, z (of Linguistics, INL). The letters c, q, w, and y are not used in Tetun, except for the proper names and international symbols. Accented vowels á, é, í, ó, ú, are also used, and the apostrophe (’) denotes a glottal stop. Additionally, the hyphen is also used to indicate mono-semantic compound words. Some basic phrases in Tetun are presented in Table 1.
@@ 3.3. Morphology
Morphological processes such as circumfixes and reduplication are also contributing to both the formation of new words and the modification of existing word structures. Circumfixes involve the simultaneous addition of a prefix and a suffix to a base word, while reduplication is a morphological process in which a part of a word is copied, either fully or partially, to form a new word that may have additional morphemes attached to it (Greksáková, 2018; Urbanczyk, 2017).
Tetun does not have rich inflectional and derivational morphology, with only a few inflectional affixes (Greksáková, 2018; Hull and Correia, 2005; van Klinken et al., 2002). Tetun affixes include both native Tetun elements and those derived from Portuguese. Prefixes are exclusive of native Tetun, whereas suffixes can derive from either native Tetun or Portuguese. In compounding, words are combined using hyphens, exclusively with native Tetun words. Additionally, Tetun utilizes circumfixes and reduplication within its native vocabulary and adopts zero derivation for Portuguese-derived words. Examples of Tetun inflection and derivation are provided in Table 2.
@@ 3.4. Portuguese Loanwords
Klinken and Hajek (van Klinken and Hajek, 2018) studied a selection of seven articles from different newspapers in 2009 and reported an average of 32% words are Portuguese loanwords. Similarly, Greksáková (Greksáková, 2018) highlighted 35% of Portuguese loanwords in the analysis of 73,892 words from interview transcripts. Moreover, Hajek and Klinken (Hajek and van Klinken, 2019) described Tetun’s influence from Portuguese in newspaper and technical writing rising to over 40%, with headlines often almost entirely in Portuguese. In a recent study, de Jesus and Nunes (de Jesus and Nunes, 2024a) reported 28.20% of Portuguese loanwords in Tetun when analyzing approximately 10.69 million words extracted from the Labadain-30k+ dataset (de Jesus and Nunes, 2024b), for an interval time from 2017 to 2023. Additionally, they observed a 5.09 percentage point increase in Portuguese loanwords when comparing documents created before and after 2017 in the Labadain-30k+ dataset. Examples of Portuguese loanwords are presented in Table 3.
@@ 4. Dataset
In this work, we used the Labadain-30k+ dataset (de Jesus and Nunes, 2024b), which consists of 33,550 Tetun documents acquired through web crawling and augmented with Tetun Wikipedia documents released by the Wikimedia Foundation (Wikimedia, 2023). The dataset was thoroughly audited by native Tetun speakers at the document level and comprised a diverse range of categories, including news articles, Wikipedia entries, legal and government documents, research papers, among others (de Jesus and Nunes, 2024a). A detailed description of the dataset is provided in Table 4, with a summary grouped by category of documents presented in Table 5. This dataset was utilized to develop the Labadain-StopwordsLabadain-Stemmer, and Labadain-Avaliadór, which were subsequently used to evaluate the retrieval effectiveness of Tetun text ad-hoc retrieval.
@@
5. Phases of Tetun Text Ad-Hoc Retrieval Development
To establish a baseline for Tetun text ad-hoc retrieval, we follow the development phases illustrated in Figure 1. The process begins with the creation of a Tetun stopword list, followed by the development of a Tetun stemmer and a test collection, and concludes with experiments to establish the baseline. Each phase is detailed in the following subsections.
Figure 1. Phases of Tetun Text Ad-Hoc Retrieval Development.
[…]
Phases of Tetun Text Ad-Hoc Retrieval Development.
@@ 5.1. Labadain-Stopwords Construction
This initial phase focuses on constructing a Tetun stopword list. Since manually creating stopword lists is both time-consuming and expensive, we adopted a corpus-based approach using the Labadain-30k+ dataset. Candidate stopwords were generated using frequency- and network-based detection methods and the resulting lists were merged and validated by two native Tetun speakers to produce the final Tetun stopword list (named Labadain-Stopwords).
Building on the findings of Ali et al. (Ali et al., 2024), which demonstrated the superior effectiveness of network-based methods compared to traditional frequency-based techniques for stopword detection, we extended the evaluation to Tetun using the Labadain-Stopwords. The results were then compared with those for Portuguese and English to gain further insight.
@@ 5.2. Labadain-Stemmer Development
This phase focused on developing the stemmer algorithms for Tetun, called Labadain-Stemmer. Since a substantial portion of Tetun verbs, nouns, and adjectives are Portuguese loanwords, and Tetun suffixes encompass Portuguese-derived words and native Tetun, we created three stemmer variants: light, moderate, and heavy. The light variant targets only the suffixes of Portuguese loanwords, the moderate variant addresses both Portuguese loanword and native Tetun suffixes, and the heavy variant handles Portuguese loanword suffixes, as well as native Tetun prefixes and suffixes.
To evaluate the proposed stemmer, we conducted both intrinsic and extrinsic assessments. For the intrinsic evaluation, we systematically extracted a subset of vocabularies from the Labadain-30k+ dataset (de Jesus and Nunes, 2024b) and collaborated with native Tetun-speaking students to construct sample words. These students assessed this list to establish a ground truth list, with each word paired with its corresponding lemma (root). The ground truth set was then used to evaluate the accuracy of each stemmer variant using Paice metrics (Paice, 1994). For extrinsic evaluation, we tested the effectiveness of Tetun stemmers in the ad-hoc retrieval task.
@@ 5.3. Labadain-Avaliadór Building
Given the absence of a test collection for Tetun text ad-hoc retrieval, this phase focused on creating a collection following TREC guidelines. Native Tetun-speaking students developed queries by examining real-world search logs and the document collection derived from the Labadain-30k+ dataset (de Jesus and Nunes, 2024b). These students also assessed the relevance of query-document pairs using a user-friendly interface we developed, with the document pooling process automated to streamline the assessment workflow. The resulting test collection for Tetun is called the Labadain-Avaliadór.
@@ 5.4. Tetun Ad-Hoc Retrieval Baseline Establishment
This phase focused on exploring various retrieval strategies for Tetun text ad-hoc retrieval. Documents and queries were initially preprocessed by converting text to lowercase, normalizing apostrophes, removing punctuation and special characters, tokenizing into individual tokens, and then performing document indexing, retrieval, and ranking to establish the baseline. Additional preprocessing steps, such as handling accented letters, apostrophes, and hyphens, were applied individually to assess their impact on retrieval effectiveness relative to the baseline. The process also included stopword removal and stemming.
The features that demonstrated improvements over the baseline were selected and combined for subsequent experiments to create the baselines. This approach was applied to both document titles (short text) and content (long text), employing various retrieval and ranking models. The effectiveness of these preprocessing steps and models was then assessed using various evaluation metrics to identify the most effective retrieval strategy for Tetun text ad-hoc retrieval.
@@ 6. Labadain-Stopwords Construction
This section presents the frequency- and network-based approaches used to develop Labadain-Stopwords, a Tetun stopword list. It describes the methodology for constructing the list and compares the effectiveness of network-based methods with traditional frequency-based techniques for stopword detection.
@@ 6.1. Overview
Network-based methods leveraged topological properties of co-occurrence networks, such as in-degree, out-degree, and degree. The in-degree represents the number of incoming connections, reflecting how often a word is preceded by others. The out-degree measures outgoing connections, showing how frequently a word leads to subsequent terms. The degree is calculated as the sum of in-degree and out-degree.
@@ 6.2. Approach
The Labadain-Stopwords was constructed using the Labadain-30k+ dataset (de Jesus and Nunes, 2024b). The process began with preprocessing steps, including lowercase, normalizing apostrophes, removing punctuation, special characters, numbers, and extra spaces, followed by tokenization with the Tetun tokenizer (de Jesus and Nunes, 2024c) and deduplication to create a vocabulary. The traditional term-weighting techniques (TF, IDF, and TF-IDF) were then applied to the vocabulary to assign weights to each word.
To analyze network properties, a vocabulary-level co-occurrence network was built as a directed graph from the preprocessed text, where each word was represented as a node. The graph properties were analyzed by calculating in-degree (total incoming links), out-degree (total outgoing links), and degree (sum of incoming and outgoing links) for each node, reflecting the connectivity of each word within the network.
Using these scores, the top 1,000 words from each method were selected in descending order based on their scores to create lists of potential stopwords. These lists were then merged, with duplicates and misspelled words excluded, to produce a candidate stopword list. Two native Tetun speakers—a Ph.D. student and an undergraduate student—reviewed and validated this list, resulting in Labadain-Stopwords, containing 160 Tetun stopwords. The complete list with English translations is provided in Appendix 12.1.
@@ 6.3. Experiment and Evaluation
In the network-based approach for detecting stopwords presented by Ali et al. (Ali et al., 2024), Tetun stopwords were manually translated from the English stopwords in the NLTK library [^7] https://www.nltk.org to establish a ground truth. In this study, we utilized the Labadain-30k+ dataset (de Jesus and Nunes, 2024b) and evaluated the effectiveness of the network-based approach using the Labadain-Stopwords. To further assess the methods’ performance across both low- and high-resource languages, we also conducted experiments with English and Portuguese.
For the Portuguese and English datasets, we used documents extracted from the CC-100 dataset (Wenzek et al., 2020) and employed stopword lists from the NLTK library as ground truth. The process of assigning weights to Portuguese and English words followed the same approach used for Tetun. A summary of the datasets used to create the stopword lists is provided in Table 6.
@@
For evaluation, we used precision at $n$ (P@$n$) to measure the proportion of stopwords among the top-$n$ words. While Ali et al. (Ali et al., 2024) limited their analysis to P@200, we extended the P@$n$ cutoff to 1,000. For this purpose, we applied intervals of approximately 25 for cutoffs up to 100, and intervals of 250 for cutoffs between 100 and 1,000.
@@ 6.4. Results
The results of the experiment with Tetun, presented in Table 7, show that network-based approaches generally outperform traditional term weighting methods in identifying stopwords. Specifically, in-degree consistently demonstrates superior performance across most cutoffs, except at P@75, where degree slightly surpasses it. At P@10, P@25, and P@1000, in-degree and degree achieve identical performance scores. Notably, at the P@10 cutoff, all techniques perform equally well, achieving perfect precision. Among traditional term weighting methods, the results are comparable, with IDF slightly outperforming TF and TF-IDF at P@500 and P@1000.
@@
In-degree 1,$0$ 0,9600 0,8400 0,7200 0,7000 0,4720 0,3080 0,2347 0,1930
Out-degree 1,$0$ 0,8800 0,8000 0,6933 0,6000 0,4240 0,2900 0,2160 0,1780
Degree 1,$0$ 0,9600 0,8200 0,7333 0,6500 0,4640 0,3000 0,2253 0,1930
TF 1,$0$ 0,9200 0,6800 0,6000 0,5200 0,3600 0,2480 0,1933 0,1610
IDF 1,$0$ 0,9200 0,6800 0,6000 0,5200 0,3600 0,2540 0,1973 0,1640
TF-IDF 1,$0$ 0,9200 0,6800 0,5867 0,5100 0,3560 0,2500 0,1947 0,1620
In the case of the Portuguese dataset, similar patterns were observed, as presented in Table 8, network-based methods again demonstrated superior performance. The degree slightly outperformed in-degree at P@50, P@250, P@500, and P@750. At P@10, P@25, P@100, and P@1000, both in-degree and out-degree achieved identical scores. In-degree outperformed degree at P@75. At P@10 and P@25, all techniques performed equally well, achieving perfect precision. At P@1000, all methods achieved identical performance scores. As for traditional term weighting approaches, they produced identical results across all evaluated cutoffs (when the dataset size was reduced, scores varied slightly at different cutoffs).
@@
In-degree 1,$0$ 1,$0$ 0,9000 0,7600 0,6700 0,3800 0,2260 0,1587 0,1260
Out-degree 1,$0$ 0,9600 0,9600 0,8000 0,6700 0,3560 0,2180 0,1573 0,1270
Degree 1,$0$ 1,$0$ 0,9400 0,8133 0,7100 0,3680 0,2240 0,1587 0,1240
TF 1,$0$ 0,9600 0,9000 0,7600 0,6500 0,3320 0,2000 0,1480 0,1200
IDF 1,$0$ 0,9600 0,9000 0,7600 0,6500 0,3320 0,2000 0,1480 0,1200
TF-IDF 1,$0$ 0,9600 0,9000 0,7600 0,6500 0,3320 0,2000 0,1480 0,1200
Similarly, in the English dataset, network-based approaches maintained their advantage, as shown in Table 9. The results mirrored those of the Portuguese dataset, with degree slightly outperforming in-degree at P@50, P@250, P@500, and P@750. In-degree and degree achieved identical scores at P@100, while in-degree outperformed degree at P@75. As with the other datasets, all methods achieved perfect precision at P@10 and P@25, and identical performance scores at P@1000. Similarly to the Portuguese dataset, the traditional term weighting methods in the English dataset yielded identical results across all evaluated cutoffs.
@@
In-degree 1,$0$ 1,$0$ 0,9400 0,9333 0,8200 0,4120 0,2280 0,1573 0,1220
Out-degree 1,$0$ 1,$0$ 0,9800 0,8400 0,7600 0,4160 0,2280 0,1573 0,1220
Degree 1,$0$ 1,$0$ 0,9800 0,9067 0,8200 0,4280 0,2300 0,1613 0,1220
TF 1,$0$ 1,$0$ 0,9600 0,8933 0,7500 0,4240 0,2280 0,1573 0,1220
IDF 1,$0$ 1,$0$ 0,9600 0,8933 0,7500 0,4240 0,2280 0,1573 0,1220
TF-IDF 1,$0$ 1,$0$ 0,9600 0,8933 0,7500 0,4240 0,2280 0,1573 0,1220
6.5. Discussions
To analyze the stopword detection approaches at the different levels, we divide precision in lower cutoffs (up to P@25), mid-range cutoffs (P@50 to P@100), and higher cutoffs (P@250 to P@750). At lower cutoffs, all methods yielded similar results, with network-based approaches, such as in-degree and degree, slightly outperforming traditional term-weighting methods by a small margin of +0.04 points in Tetun and Portuguese at P@25.

### v6

@@ 2.4. Summary
In ad-hoc text retrieval, preprocessing steps, such as stopword removal and stemming, are often employed to enhance retrieval efficiency and effectiveness. However, studies show that the impact of these techniques on retrieval effectiveness varies across languages, proving beneficial in some instances but less so in others. Furthermore, Ghosh and Bhattacharya (2017) highlighted that the influence of stopwords can differ even within the same language across different collections, such as Bangla and Hindi in the FIRE datasets of 2010 and 2011.
@@
This study addresses a critical gap in Tetun text retrieval by introducing three essential resources: a stopword list (Labadain-Stopwords), a language-specific stemmer (Labadain-Stemmer), and a Tetun test collection (Labadain-Avaliadór). Through a series of experiments, we evaluate various retrieval strategies to establish baselines and identify the most effective approach for ad-hoc text retrieval in Tetun. The subsequent sections provide a detailed overview of the development of each resource and its application in the experiments.
@@ 3. Tetun
This section presents an overview of Tetun, including its orthography, morphology, and Portuguese loanwords.
@@ 3.1. Overview
Tetun, alternatively written as Tetum (in English) or Tétum (in Portuguese), is an Austronesian language spoken in Timor-Leste, an island nation in Southeast Asia. It has two primary varieties: Tetun Dili, also known as Tetun Prasa (commonly referred to simply as Tetun), and Tetun Terik (van Klinken et al., 2002). Tetun has two standardized forms: one developed by the Instituto Nacional de Linguística (Tetun INL) and another by the Dili Institute of Technology (Tetun DIT). Tetun Terik, meanwhile, remains one of the dialects spoken in Timor-Leste.
Tetun is one of Timor-Leste’s official languages alongside Portuguese (Vasconcelos et al., 2011). The government recognized Tetun INL as the official Tetun, which is used in the education system, official publications, and media (DL 01/2004, 2004). Tetun DIT was developed by linguists at the Dili Institute of Technology with some standardized differences from Tetun INL in terms of writing conventions (van Klinken et al., 2002). For example, the words [  (give), ne’ebé (which/that) ] in Tetun INL correspond to [ foo, neebe ] in Tetun DIT. According to the 2015 census report, Timor-Leste’s population was 1.18 million, with 78.78% of the population being Tetun speakers (de Jesus and Nunes, 2024a). Among them, 30.50% considered Tetun their home language, while 48.28% spoke it as a second or third language. Census 2022 reported a population growth of 13.40%, increasing from 1.18 million to 1.34 million (INETL, 2022), but it did not provide specific indicators for Tetun speakers.
@@ 3.2. Orthography
Tetun INL is based on the Latin alphabet, consisting of 5 vowels: a, e, i, o, u, and 21 consonants: b, d, f, g, h, j, k, l, ll, m, n, ñ, p, r, rr, s, t, ’, v, x, z (of Linguistics, INL). The letters c, q, w and y are not used in Tetun except for the proper names and international symbols. Accented vowels á, é, í, ó, ú are also used, and the apostrophe (’) denotes a glottal stop. Additionally, the hyphen is also used to indicate monosemantic compound words. Some basic phrases in Tetun are presented in Table 1.
@@ 3.3. Morphology
Morphological processes such as circumfixes and reduplication also contribute to both the formation of new words and the modification of existing word structures. Circumfixes involve the simultaneous addition of a prefix and a suffix to a base word, while reduplication is a morphological process in which a part of a word is copied, either fully or partially, to form a new word that may have additional morphemes attached to it (Greksáková, 2018; Urbanczyk, 2017).
Tetun does not have rich inflectional and derivational morphology, with only a few inflectional affixes (Greksáková, 2018; Hull and Correia, 2005; van Klinken et al., 2002). Tetun affixes include both native Tetun elements and those derived from Portuguese. Prefixes are exclusive of native Tetun, whereas suffixes can derive from either native Tetun or Portuguese. In compounding, words are combined using hyphens, exclusively with native Tetun words. Additionally, Tetun uses circumfixes and reduplication within its native vocabulary and adopts zero derivation for Portuguese-derived words. Examples of Tetun inflection and derivation are provided in Table 2.
@@ 3.4. Portuguese Loanwords
Klinken and Hajek (van Klinken and Hajek, 2018) studied a selection of seven articles from different newspapers in 2009 and reported an average of 32% of words are Portuguese loanwords. Similarly, Greksáková (2018) highlighted 35% of Portuguese loanwords in the analysis of 73,892 words from interview transcripts. Moreover, Hajek and Klinken (Hajek and van Klinken, 2019) described Tetun’s influence from Portuguese in newspaper and technical writing, rising to over 40%, with headlines often almost entirely in Portuguese. In a recent study, de Jesus and Nunes (2024a) reported 28.20% of Portuguese loanwords in Tetun when analyzing approximately 10.69 million words extracted from the Labadain-30k+ dataset (de Jesus and Nunes, 2024b) for an interval time from 2017 to 2023. Additionally, they observed a 5.09 percentage point increase in Portuguese loanwords when comparing documents created before and after 2017 in the Labadain-30k+ dataset. Examples of Portuguese loanwords are presented in Table 3.
@@ 4. Dataset
In this work, we employed the Labadain-30k+ dataset (de Jesus and Nunes, 2024b), comprising 33,550 Tetun documents acquired through web crawling. The dataset was thoroughly audited by native Tetun speakers at the document level and comprised a diverse range of categories, including news articles, Wikipedia entries, legal and government documents, and research papers, among others (de Jesus and Nunes, 2024a). A detailed description of the dataset is provided in Table 4, with a summary grouped by category of documents presented in Table 5. This dataset was employed to develop Labadain-Stopwords, Labadain-Stemmer, and Labadain-Avaliadór, which were subsequently used to evaluate the retrieval effectiveness of Tetun ad-hoc text retrieval.
@@
5. Methodology
To establish baselines for Tetun ad-hoc text retrieval, we employ the methodology illustrated in Figure 1. The process begins with the creation of a Tetun stopword list, continues with the development of a stemmer and a test collection, and concludes with experiments to establish the baselines. Each stage is described in the following subsections.
Figure 1. Methodology for Establishing Baselines in Tetun Ad-Hoc Text Retrieval.:
@@ 5.1. Labadain-Stopwords Construction
This initial stage focuses on constructing a Tetun stopword list. Since manually creating stopword lists is both time-consuming and expensive, we adopted a corpus-based approach using the Labadain-30k+ dataset. Candidate stopwords were generated using frequency- and network-based detection methods, and the resulting lists were merged and validated by two native Tetun speakers to produce the final Tetun stopword list (called Labadain-Stopwords).
Building on the findings of Ali et al. (2024), which demonstrated the superior effectiveness of network-based methods compared to traditional frequency-based techniques for stopword detection, we extended the evaluation to Tetun using the Labadain-Stopwords as the ground-truth set. We then compared the results with those for Portuguese and English to gain further insight.
@@ 5.2. Labadain-Stemmer Development
This stage focused on developing the stemmer algorithms for Tetun, called Labadain-Stemmer. Since a substantial portion of Tetun verbs, nouns, and adjectives are Portuguese loanwords, and Tetun suffixes encompass Portuguese-derived words and native Tetun, we created three stemmer variants: light, moderate, and heavy. The light variant removes only the suffixes of Portuguese loanwords, the moderate variant addresses both Portuguese loanwords and native Tetun suffixes, and the heavy variant handles Portuguese loanword suffixes, as well as native Tetun prefixes and suffixes.
To evaluate the proposed stemmer, we conducted both intrinsic and extrinsic assessments. For the intrinsic evaluation, we systematically extracted a subset of vocabularies from the Labadain-30k+ dataset (de Jesus and Nunes, 2024b) and collaborated with native Tetun-speaking students to construct a sample of words. These students assessed the sample word list provided to establish a ground truth list, with each word paired with its corresponding lemma (root). The ground truth set was then used to evaluate the accuracy of each stemmer variant using the Paice metrics (Paice, 1994). For extrinsic evaluation, we tested the effectiveness of Tetun stemmers in the ad-hoc text retrieval task.
@@ 5.3. Labadain-Avaliadór Building
Since no test collection exists for Tetun ad-hoc text retrieval, this stage focused on creating one following TREC guidelines. Native Tetun-speaking students developed queries by examining real-world search logs and the document collection sourced from the Labadain-30k+ dataset (de Jesus and Nunes, 2024b). The same students also assessed the relevance of query-document pairs using a user-friendly interface we developed, with the document pooling process automated to streamline the assessment workflow. The resulting Tetun test collection is called Labadain-Avaliadór.
@@ 5.4. Tetun Ad-Hoc Retrieval Baseline Establishment
This stage focused on investigating various retrieval strategies for Tetun ad-hoc text retrieval. Documents and queries were initially preprocessed by converting text to lowercase, normalizing apostrophes, removing punctuation and special characters, tokenizing into individual tokens, and then performing document indexing, retrieval, and ranking to establish the baselines. Additional preprocessing steps, such as handling accented letters, apostrophes, and hyphens, were applied individually to assess their impact on retrieval effectiveness relative to the baselines. The process also included stopword removal and stemming.
The features that demonstrated improvements over the baseline were selected and combined for subsequent experiments to create the baselines. This approach was applied to both document titles (short text) and content (long text), employing various retrieval and ranking models. The effectiveness of these preprocessing steps and models was then assessed using various evaluation metrics to identify the most effective retrieval strategy for Tetun ad-hoc text retrieval.
@@ 6. Labadain-Stopwords Construction
This section introduces the frequency- and network-based approaches used to create Labadain-Stopwords, a Tetun stopword list. It describes the methodology for constructing the list and compares the effectiveness of network-based methods with traditional frequency-based techniques for stopword detection.
@@ 6.1. Overview
Network-based methods exploit the topological properties of co-occurrence networks modeled as directed graphs, including in-degree, out-degree, and degree. The in-degree represents the number of incoming connections, indicating how often a word is preceded by others. The out-degree captures the number of outgoing connections, showing how frequently a word precedes subsequent terms. The degree is defined as the sum of the in-degree and out-degree.
@@ 6.2. Approach
The Labadain-Stopwords was constructed using the Labadain-30k+ dataset (de Jesus and Nunes, 2024b). The process began with preprocessing steps, including lowercase, normalizing apostrophes, removing punctuation, special characters, numbers, and extra spaces, followed by tokenization using the Tetun tokenizer (de Jesus and Nunes, 2024e) and deduplication to create a vocabulary. The traditional term-weighting techniques (TF, IDF, and TF-IDF) were then applied to the vocabulary to assign weights to each word.
To analyze network properties, we constructed a vocabulary-level co-occurrence network as a directed graph from the preprocessed text, where each word corresponds to a node. For each node, we calculated in-degree (number of incoming links), out-degree (number of outgoing links), and degree (the sum of incoming and outgoing links), thereby quantifying the connectivity of words within the network.
Using these scores, the top 1,000 words from each method were selected in descending order based on their scores to create lists of potential stopwords. These lists were then merged, with duplicates and misspelled words excluded, to produce a candidate stopword list. Two native Tetun speakers—a Ph.D. student and an undergraduate student—reviewed and validated this list, resulting in Labadain-Stopwords, containing 160 Tetun stopwords (de Jesus and Nunes, 2025c). The complete list with English translations is provided in Appendix 12.1.
@@ 6.3. Experiment and Evaluation
In the network-based approach for detecting stopwords proposed by Ali et al. (2024), Tetun stopwords were manually translated from the English stopwords in NLTK [^7] https://www.nltk.org to establish the ground-truth set. In this study, we used the Labadain-30k+ dataset (de Jesus and Nunes, 2024b) and evaluated the effectiveness of the approach with Labadain-Stopwords. To further assess performance across both low- and high-resource languages, we also conducted experiments with English and Portuguese.
For Portuguese and English, we used documents extracted from the CC-100 dataset (Wenzek et al., 2020) and employed stopword lists from NLTK as the ground truth. The process of assigning weights to Portuguese and English words followed the same approach used for Tetun. A summary of the datasets used to create the stopword lists is provided in Table 6.
@@
For evaluation, we used precision at $n$ (P@$n$) to measure the proportion of stopwords among the top-$n$ words. While Ali et al. (2024) limited their analysis to P@200, we extended the P@$n$ cutoff to 1,000. For this purpose, we applied intervals of approximately 25 for cutoffs up to 100 and intervals of 250 for cutoffs between 100 and 1,000.
@@ 6.4. Results
The results of the experiment with Tetun are presented in Table 7, demonstrating that network-based approaches generally outperform traditional term weighting methods in identifying stopwords. Specifically, in-degree consistently demonstrates superior performance across most cutoffs, except at P@75, where degree slightly surpasses it. At P@10, P@25, and P@1000, in-degree and degree achieve identical performance scores. Notably, at the P@10 cutoff, all techniques perform equally well, achieving perfect precision. Among traditional term weighting methods, the results are comparable, with IDF slightly outperforming TF and TF-IDF at P@500 and P@1000.
@@
In-degree 1.0000 0.9600 0.8400 0.7200 0.7000 0.4720 0.3080 0.2347 0.1930
Out-degree 1.0000 0.8800 0.8000 0.6933 0.6000 0.4240 0.2900 0.2160 0.1780
Degree 1.0000 0.9600 0.8200 0.7333 0.6500 0.4640 0.3000 0.2253 0.1930
TF 1.0000 0.9200 0.6800 0.6000 0.5200 0.3600 0.2480 0.1933 0.1610
IDF 1.0000 0.9200 0.6800 0.6000 0.5200 0.3600 0.2540 0.1973 0.1640
TF-IDF 1.0000 0.9200 0.6800 0.5867 0.5100 0.3560 0.2500 0.1947 0.1620
When evaluated on Portuguese, similar patterns were observed, as shown in Table 8, with network-based methods again demonstrating superior performance. The degree slightly surpassed the in-degree at P@50, P@250, P@500, and P@750. At P@10, P@25, P@100, and P@1000, both in-degree and out-degree achieved identical scores. In-degree outperformed degree at P@75. At P@10 and P@25, all techniques performed equally well, achieving perfect precision. At P@1000, all methods achieved identical scores. Traditional term-weighting approaches yielded identical results across all evaluated cutoffs, though slight variations appeared at certain cutoffs when the dataset size was reduced.
@@
In-degree 1.0000 1.0000 0.9000 0.7600 0.6700 0.3800 0.2260 0.1587 0.1260
Out-degree 1.0000 0.9600 0.9600 0.8000 0.6700 0.3560 0.2180 0.1573 0.1270
Degree 1.0000 1.0000 0.9400 0.8133 0.7100 0.3680 0.2240 0.1587 0.1240
TF 1.0000 0.9600 0.9000 0.7600 0.6500 0.3320 0.2000 0.1480 0.1200
IDF 1.0000 0.9600 0.9000 0.7600 0.6500 0.3320 0.2000 0.1480 0.1200
TF-IDF 1.0000 0.9600 0.9000 0.7600 0.6500 0.3320 0.2000 0.1480 0.1200
Similarly, in English, network-based approaches maintained their advantage, as shown in Table 9. The results mirrored those of Portuguese, with degree slightly outperforming in-degree at P@50, P@250, P@500, and P@750. In-degree and degree attained identical scores at P@100, while in-degree outperformed degree at P@75. As with the other languages, all methods achieved perfect precision at P@10 and P@25 and identical scores at P@1000. As in Portuguese, traditional term-weighting methods in English yielded identical results across all evaluated cutoffs.
@@
In-degree 1.0000 1.0000 0.9400 0.9333 0.8200 0.4120 0.2280 0.1573 0.1220
Out-degree 1.0000 1.0000 0.9800 0.8400 0.7600 0.4160 0.2280 0.1573 0.1220
Degree 1.0000 1.0000 0.9800 0.9067 0.8200 0.4280 0.2300 0.1613 0.1220
TF 1.0000 1.0000 0.9600 0.8933 0.7500 0.4240 0.2280 0.1573 0.1220
IDF 1.0000 1.0000 0.9600 0.8933 0.7500 0.4240 0.2280 0.1573 0.1220
TF-IDF 1.0000 1.0000 0.9600 0.8933 0.7500 0.4240 0.2280 0.1573 0.1220
6.5. Discussion
To examine the stopword detection approaches across different levels, we divide precision into lower cutoffs (up to P@25), mid-range cutoffs (P@50 to P@100), and higher cutoffs (P@250 to P@750). At lower cutoffs, all methods yielded similar results, with network-based approaches, such as in-degree and degree, slightly outperforming traditional term-weighting methods by a small margin of +0.04 points in Tetun and Portuguese at P@25.